OpenStudy (anonymous):

Corey needs 24 boards that are 47 1/2 inches long. How many feet of boards should he buy? Explain. If you can only buy 8 foot boards , how many should he buy??

OpenStudy (31356):

1. (24 * 47.5) / 12 = ? 2. 8 * 12 = 96 inches 96 / 47.5 = 2.02 He can get two 47.50-inch lengths out of each 8-foot board. So he needs 12-foot boards.

OpenStudy (31356):

Medal?

OpenStudy (31356):

Thanks!